OpenStudy (anonymous):

a close box with a square base is to have a vlume of 1000cu.m. the material for the top and bottom of the box is to cost P100/sq.m. and the material for the sides is to cost twice the top. find the dimensions of the box so that the total cost of the material is minimum..?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

say the side of the square base is \(\large x\) so the area of bottom = area of top = \(\large ?\)

OpenStudy (anonymous):

right..

OpenStudy (anonymous):

x^2

ganeshie8 (ganeshie8):

yes and since `the material for the top and bottom of the box is to cost P100/sq.m. ` cost of top and bottom = \(\large (2x^2)\times 100\)

OpenStudy (anonymous):

therefore it becomes 200x^2 right..?

ganeshie8 (ganeshie8):

yes thats the expression for cost of top and bottom, lets work the lateral sides

ganeshie8 (ganeshie8):

say the height of box is \(h\) since the box has four lateral sides, area of sides = \(4(x\times h) = 4xh \)

ganeshie8 (ganeshie8):

lets work again from this step

OpenStudy (anonymous):

by the way why do we set the derivative equal to zero before?

ganeshie8 (ganeshie8):

` and the material for the sides is to cost twice the top.` that means the cost of sides = \(\large 4xh \times 200 = 800xh\)

ganeshie8 (ganeshie8):

So the total cost would be : \(200x^2 + 800xh\)

ganeshie8 (ganeshie8):

see the mistake we committed earlier ?

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

use the volume information and eliminate \(h\) like we did earlier : \(x^2h = 1000 \implies h = \dfrac{1000}{x^2} \)

ganeshie8 (ganeshie8):

cost \(200x^2 + 800xh = 200x^2 + 800x \dfrac{1000}{x^2}\) \(= 200x^2 + \dfrac{800000}{x} \)

ganeshie8 (ganeshie8):

differentiate this function set it equal to 0 and solve x

OpenStudy (anonymous):

back to my question earlier why did we equate it to zero...?

OpenStudy (anonymous):

nevermind that question.. i think its the rule to get the value of x... any way thanks a lot.. you really are great....

ganeshie8 (ganeshie8):

nope, thats a very good question actually. lets see why setting the derivative equal to 0 works

ganeshie8 (ganeshie8):

you familiar with parabolas , right ?

ganeshie8 (ganeshie8):